# The Limits of Assumption-free Tests for Algorithm Performance

### v1

@@ A.2 Proof of Theorem 1
where $\mathcal{D}_{N}$ denotes an i.i.d. sample from $P$, and $\mathcal{D}^{(1)},\dots,\mathcal{D}^{({\widehat{r}})}$ denote the data sets constructed when running ${\widehat{T}}(\mathcal{A},\mathcal{D}_{N})$. Let $\mathcal{E}_{*}$ be the event that $(x_{*},y_{*})\not\in\mathcal{D}_{N}\cup\mathcal{D}^{(1)}\cup\dots\cup\mathcal{D}^{({\widehat{r}})}$, so that (*$\mathbb{P}(\mathcal{E}_{*})\geq 1-\epsilon$.
@@
Next, observe that on the event $\mathcal{E}_{*}$, we have ${\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{N})={\widehat{T}}(\mathcal{A},\mathcal{D}_{N})$, almost surely. Informally, this is because when running the black-box text ${\widehat{T}}$, if we initialize at the same data set $\mathcal{D}_{N}$, and we never observe the data point $(x_{*},y_{*})$, then $\mathcal{A}$ and $\mathcal{A}^{\prime}$ are returning the same fitted models at every iteration, and so all iterations of the tests ${\widehat{T}}(\mathcal{A},\mathcal{D}_{N})$ and ${\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{N})$ are identical to each other. (A more formal proof of this type of statement can be found in the proof of Kim and Barber, (2023, Theorem 2).) In particular, we have
@@
where the last step holds by definition of $C$. Therefore, by validity of ${\widehat{T}}$ (5), we must have
@@
A.3 Proof of Theorem 2
A.4 Proof of Theorem 4
@@
A.4.1 Step 1: some preliminaries
First, if $1-c<(1/\alpha)^{-1/N}$, then the desired bound (21) holds trivially, since the right-hand side of (21) is $>1$ in this case. Therefore we can restrict our attention to the nontrivial case where $1-c\geq(1/\alpha)^{-1/N}$. This implies that
@@ A.4.1 Step 1: some preliminaries
A.4.2 Step 2: constructing a new algorithm and a new distribution
We begin by explaining the intuition behind the construction of $P^{\prime}$ and $\mathcal{A}^{\prime}$, which is motivated by the procedure of the black-box test. As in the proof of Theorem 1, we want $\mathcal{A}^{\prime}$ to behave in the same way as $\mathcal{A}$ on a typical dataset drawn from $P$, so that ${\widehat{T}}$ cannot easily distinguish between $\mathcal{A}$ and $\mathcal{A}^{\prime}$—but on the other hand, we want $R_{P,n}(\mathcal{A}^{\prime})\geq\tau$ so that ${\widehat{T}}$ is constrained to have a low probability of returning a $1$ when run on $\mathcal{A}^{\prime}$ (and consequently, also when run on the original algorithm of interest $\mathcal{A}$).
@@ A.4.2 Step 2: constructing a new algorithm and a new distribution
We will show below that the new algorithm $\mathcal{A}^{\prime}$ and distribution $P^{\prime}$ satisfy $R_{P^{\prime},n}(\mathcal{A}^{\prime})\geq\tau$ and $\beta_{q}(\mathcal{A}^{\prime},P^{\prime},n)\leq\gamma_{q}$, as long as we choose $\epsilon$ to be sufficiently small. This means that we must have $\mathbb{P}({\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{N}^{\prime})=1)\leq\alpha$, by the assumption (8) on the validity of ${\widehat{T}}$ for testing stable algorithms, and so we have
@@
A.4.3 Step 3: verifying the risk
@@
A.4.4 Step 4: verifying the stability
In this step we verify that $\beta_{q}(\mathcal{A}^{\prime},P^{\prime},n)\leq\gamma_{q}$. Let $(X_{1},Y_{1}),\dots,(X_{n+1},Y_{n+1})\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$, and let $(X_{1}^{\prime},Y_{1}^{\prime}),\dots,(X_{n+1}^{\prime},Y_{n+1}^{\prime})\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P^{\prime}$. Define $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i\in[n]}$, let $\mathcal{D}_{n+1}=\{(X_{i},Y_{i})\}_{i\in[n+1]}$, and let $\mathcal{D}_{n}^{(-j)}=\{(X_{i},Y_{i})\}_{i\in[n]\backslash\{j\}}$. Define $\mathcal{D}^{\prime}_{n}$, $\mathcal{D}^{\prime}_{n+1}$, and $\mathcal{D}^{\prime}_{n}{}^{(-j)}$ analogously. Let ${\widehat{f}}_{n}=\mathcal{A}(\mathcal{D}_{n};\xi)$ and ${\widehat{f}}_{n}^{(-j)}=\mathcal{A}(\mathcal{D}_{n}^{(-j)};\xi)$, and ${\widehat{f}}_{n}^{\prime}=\mathcal{A}^{\prime}(\mathcal{D}_{n}^{\prime};\xi)$ and ${\widehat{f}}_{n}^{\prime}{}^{(-j)}=\mathcal{A}^{\prime}(\mathcal{D}^{\prime}_{n}{}^{(-j)};\xi)$.
@@ A.4.4 Step 4: verifying the stability
where the last step holds since data points $\{(X_{i}^{\prime},Y_{i}^{\prime})\}_{i\in[n+1]}$ are exchangeable, so if $(x_{*},y_{*})$ appears at least once in a data set of size $n+1$, with probability at least $\frac{n-1}{n+1}$ it appears in the subset given by the $n-1$ many indices $[n]\backslash\{j\}$. Combining these calculations,
@@
Appendix B Proofs: hardness results for algorithm comparison
B.1 Proof of Theorem 5
Theorem 5 can be viewed as a special case of Theorem 7. Specifically, since the comparison function takes values in $[-B,B]$, it holds trivially for any $(\mathcal{A}_{0},\mathcal{A}_{1},P,n)$ that $\beta_{q}(\mathcal{A}_{0},\mathcal{A}_{1},P,n)\leq 2B$. Thus, the result of Theorem 2 is implied by applying Theorem 7 with stability parameter $\gamma_{q}=2B$.
B.2 Proof of Theorem 6
@@
B.2.1 Step 1: some preliminaries
@@
B.2.2 Step 2: constructing new algorithms and a new distribution
@@
$\mathcal{A}_{0}^{\prime}(\mathcal{D};\xi)=\begin{cases}f_{1},&\text{ if }(x^{*},y^{*})\in\mathcal{D},\\ \mathcal{A}_{0}(\mathcal{D};\xi),&\text{ if }(x^{*},y^{*})\notin\mathcal{D}.\end{cases}\quad\quad\mathcal{A}_{1}^{\prime}(\mathcal{D};\xi)=\begin{cases}f_{0},&\text{ if }(x^{*},y^{*})\in\mathcal{D},\\ \mathcal{A}_{1}(\mathcal{D};\xi),&\text{ if }(x^{*},y^{*})\notin\mathcal{D}.\end{cases}$
@@
We will show below that the new algorithms $\mathcal{A}_{0}^{\prime},\mathcal{A}_{1}$ and distribution $P^{\prime}$ satisfy $\Delta_{P^{\prime},n}(\mathcal{A}_{0}^{\prime},\mathcal{A}_{1}^{\prime})\leq 0$ and $\beta_{q}(\mathcal{A}_{l}^{\prime},P^{\prime},n)\leq\gamma_{q}$ for $l=0,1$, as long as we choose $\epsilon$ to be sufficiently small. This means that we must have $\mathbb{P}({\widehat{T}}(\mathcal{A}_{0}^{\prime},\mathcal{A}_{1}^{\prime},\mathcal{D}_{N}^{\prime})=1)\leq\alpha$, by the assumption (14) on the validity of ${\widehat{T}}$ for comparing stable algorithms, which completes the proof exactly as for Theorem 4.
B.2.3 Step 3: verifying the risk
@@
$\Delta_{P^{\prime},n}(\mathcal{A}_{0}^{\prime},\mathcal{A}_{1}^{\prime})\leq(1-c)^{n+1}\left(R_{P_{*},n}(\mathcal{A}_{0})-R_{P_{*},n}(\mathcal{A}_{1})\right)+(1-c)\left[1-(1-c)^{n}\right]\left(R_{P_{*}}(f_{1})-R_{P_{*}}(f_{0})\right)+cB\\ =(1-c)^{n+1}\Delta_{P_{*},n}(\mathcal{A}_{0},\mathcal{A}_{1})+(1-c)\left[1-(1-c)^{n}\right]\Delta_{P_{*}}(f_{1},f_{0})+cB.$
@@
B.2.4 Step 4: verifying the stability
@@ B.3 Proof of Theorem 7
~$\tilde{\ell}(({\widehat{y}}_{0},{\widehat{y}}_{1}),y)\in[0,2B],\quad\tilde{R}_{P,n}(\tilde{\mathcal{A}})=B-\Delta_{P,n}(\mathcal{A}_{0},\mathcal{A}_{1})$
@@
Plugging in these substitutions, the result of Theorem 4 tells us that power for evaluating the risk $\tilde{R}_{P,n}(\tilde{\mathcal{A}})$ is bounded as
@@
Appendix C Additional proofs
C.1 Proof of Theorem 3
@@
$\begin{split}\mathbb{P}({\widehat{T}}(\mathcal{A},\mathcal{D}_{N})=1)&=\mathbb{P}(S<k_{*})+a_{*}\mathbb{P}(S=k_{*})\\ &=(1-a_{*})\mathbb{P}(S<k_{*})+a_{*}\mathbb{P}(S\leq k_{*})\\ &=(1-a_{*})\mathbb{P}(\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},R_{P,n}(\mathcal{A}))<k_{*})+a_{*}\mathbb{P}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},R_{P,n}(\mathcal{A}))\leq k_{*})\\ &\leq(1-a_{*})\mathbb{P}(\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},\tau)<k_{*})+a_{*}\mathbb{P}(\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},\tau)\leq k_{*})\\ &=\mathbb{P}(\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},\tau)<k_{*})+a_{*}\mathbb{P}(\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},\tau)=k_{*})\\ &=\alpha\end{split}$
where the last step holds by our choice of $k_{*},a_{*}$, while the inequality holds because $R_{P,n}(\mathcal{A})\geq\tau$ and so $\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},R_{P,n}(\mathcal{A}))$ stochastically dominates $\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},\tau)$. Next, we compute the power when $\alpha<(1-\tau)^{\lfloor N/(n+1)\rfloor}$. By this bound on $\alpha$ we have $\mathbb{P}(\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor}\tau)=0)=(1-\tau)^{\lfloor N/(n+1)\rfloor}>\alpha$, which implies that
@@
Therefore, by the construction of the Binomial test ${\widehat{T}}$, we have

### v4

@@ B.3 Proof of Theorem 3
where $\mathcal{D}_{N}$ denotes an i.i.d. sample from $P$, and $\mathcal{D}^{(1)},\dots,\mathcal{D}^{({\widehat{r}})}$ denote the data sets constructed when running ${\widehat{T}}(\mathcal{A},\mathcal{D}_{N})$. Let $\mathcal{E}_{*}$ be the event that $(x_{*},y_{*})\not\in\mathcal{D}_{N}\cup\mathcal{D}^{(1)}\cup\dots\cup\mathcal{D}^{({\widehat{r}})}$, so that $\mathbb{P}(\mathcal{E}_{*})\geq 1-\epsilon$.
@@
Next, observe that on the event $\mathcal{E}_{*}$, we have ${\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{N})={\widehat{T}}(\mathcal{A},\mathcal{D}_{N})$, almost surely. Informally, this is because when running the black-box text $^{\widehat{T}}$, if we initialize at the same data set $\mathcal{D}_{N}$, and we never observe the data point $(x_{*},y_{*})$, then $\mathcal{A}$ and $\mathcal{A}^{\prime}$ are returning the same fitted models at every iteration, and so all iterations of the tests ${\widehat{T}}(\mathcal{A},\mathcal{D}_{N})$ and ${\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{N})$ are identical to each other. (A more formal proof of this type of statement can be found in the proof of Kim and Barber, (2023, Theorem 2).) In particular, we have
@@
where the last step holds by definition of $C$. Therefore, by validity of $^{\widehat{T}}$ (6), we must have
@@
B.4 Proof of Theorem 4
@@
B.4.1 Step 1: some preliminaries
First, if $1-c<(1/\alpha)^{-1/N}$, then no matter whether (26) holds or not, the desired bound (27) holds trivially, since the right-hand side of (27) is >1>1 in this case. Therefore we can restrict our attention to the nontrivial case where $1-c\geq(1/\alpha)^{-1/N}$. This implies that
@@ B.4.1 Step 1: some preliminaries
B.4.2 Step 2: constructing a new algorithm and a new distribution
We begin by explaining the intuition behind the construction of $P^{\prime}$ and $\mathcal{A}^{\prime}$, which is motivated by the procedure of the black-box test. As in the proof of Theorem 3, we want $\mathcal{A}^{\prime}$ to behave in the same way as $\mathcal{A}$ on a typical dataset drawn from $P$, so that $^{\widehat{T}}$ cannot easily distinguish between $\mathcal{A}$ and $\mathcal{A}^{\prime}$—but on the other hand, we want $R_{P,n}(\mathcal{A}^{\prime})\geq\tau$ so that $^{\widehat{T}}$ is constrained to have a low probability of returning a $1$ when run on $\mathcal{A}^{\prime}$ (and consequently, also when run on the original algorithm of interest $\mathcal{A}$).
@@ B.4.2 Step 2: constructing a new algorithm and a new distribution
We will show below that the new algorithm $\mathcal{A}^{\prime}$ and distribution $P^{\prime}$ satisfy $R_{P^{\prime},n}(\mathcal{A}^{\prime})\geq\tau$ and $\beta_{q}(\mathcal{A}^{\prime},P^{\prime},n)\leq\gamma_{q}$, as long as we choose $\epsilon$ to be sufficiently small. This means that we must have $\mathbb{P}({\widehat{T}}(\mathcal{A}^{\prime},\mathcal{D}_{N}^{\prime})=1)\leq\alpha$, by the assumption (11) on the validity of $^{\widehat{T}}$ for testing stable algorithms, and so we have
@@
B.4.3 Step 3: verifying the risk
@@
B.4.4 Step 4: verifying the stability
In this step we verify that $\beta_{q}(\mathcal{A}^{\prime},P^{\prime},n)\leq\gamma_{q}$. Let $(X_{1},Y_{1}),\dots,(X_{n+1},Y_{n+1})\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$, and let
$(X_{1}^{\prime},Y_{1}^{\prime}),\dots,(X_{n+1}^{\prime},Y_{n+1}^{\prime})\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P^{\prime}.$
[…]
@@ B.4.4 Step 4: verifying the stability
[…]
@@
Appendix C Proofs: hardness results for algorithm comparison
C.1 Proof of Theorem 5
Theorem 5 can be viewed as a special case of Theorem 7. Specifically, since the comparison function takes values in [B,B][-B,B], it holds trivially for any $(\mathcal{A}_{0},\mathcal{A}_{1},P,n)$ that $\beta_{q}(\mathcal{A}_{0},\mathcal{A}_{1},P,n)\leq 2B$. Thus, the result of Theorem 1 is implied by applying Theorem 7 with stability parameter $\gamma_{q}=2B$.
C.2 Proof of Theorem 6
@@
C.2.1 Step 1: some preliminaries
@@
C.2.2 Step 2: constructing new algorithms and a new distribution
@@
[…]
@@
We will show below that the new algorithms $\mathcal{A}_{0}^{\prime},\mathcal{A}_{1}$ and distribution $P^{\prime}$ satisfy $\Delta_{P^{\prime},n}(\mathcal{A}_{0}^{\prime},\mathcal{A}_{1}^{\prime})\leq 0$ and $\beta_{q}(\mathcal{A}_{l}^{\prime},P^{\prime},n)\leq\gamma_{q}$ for $l=0,1$, as long as we choose $\epsilon$ to be sufficiently small. This means that we must have $\mathbb{P}({\widehat{T}}(\mathcal{A}_{0}^{\prime},\mathcal{A}_{1}^{\prime},\mathcal{D}_{N}^{\prime})=1)\leq\alpha$, by the assumption (17) on the validity of $^{\widehat{T}}$ for comparing stable algorithms, which completes the proof exactly as for Theorem 4.
C.2.3 Step 3: verifying the risk
@@
$\Delta_{P^{\prime},n}(\mathcal{A}_{0}^{\prime},\mathcal{A}_{1}^{\prime})\\ \leq(1-c)^{n+1}\left(R_{P_{*},n}(\mathcal{A}_{0})-R_{P_{*},n}(\mathcal{A}_{1})\right)+(1-c)\left[1-(1-c)^{n}\right]\left(R_{P_{*}}(f_{1})-R_{P_{*}}(f_{0})\right)+cB\\ =(1-c)^{n+1}\Delta_{P_{*},n}(\mathcal{A}_{0},\mathcal{A}_{1})+(1-c)\left[1-(1-c)^{n}\right]\Delta_{P_{*}}(f_{1},f_{0})+cB.$
@@
C.2.4 Step 4: verifying the stability
@@ C.3 Proof of Theorem 7
$\tilde{\ell}(({\widehat{y}}_{0},{\widehat{y}}_{1}),y)\in[0,2B],\quad\tilde{R}_{P,n}(\tilde{\mathcal{A}})=B-\Delta_{P,n}(\mathcal{A}_{0},\mathcal{A}_{1})$
@@
[…]
@@
Appendix D Additional proofs
D.1 Proof of Proposition 1
Let us first introduce another testing problem so that we can connect the testing problem (7) with the original testing problem in (1):
$\tilde{H}_{0}:R_{P,n}(\mathcal{A})\geq\delta B+\epsilon\quad\textnormal{versus}\quad\tilde{H}_{1}:R_{P,n}(\mathcal{A})<\delta B+\epsilon.$
Then we observe that for any $\mathcal{A},P$ such that $R_{P,n}(\mathcal{A})\geq\delta B+\epsilon$, we also have $\mathbb{P}(R_{P}(\mathcal{A}(\mathcal{D}_{n};\xi))\geq\epsilon)\geq\delta$. This is because if $\mathbb{P}(R_{P}(\mathcal{A}(\mathcal{D}_{n};\xi))\geq\epsilon)<\delta$, then
$\begin{split}R_{P,n}(\mathcal{A})&=\mathbb{E}[R_{P}(\mathcal{A}(\mathcal{D}_{n};\xi)){\mathds{1}}\{R_{P}(\mathcal{A}(\mathcal{D}_{n};\xi))\geq\epsilon\}]+\mathbb{E}[R_{P}(\mathcal{A}(\mathcal{D}_{n};\xi)){\mathds{1}}\{R_{P}(\mathcal{A}(\mathcal{D}_{n};\xi))<\epsilon\}]\\ &\leq B\mathbb{P}(R_{P}(\mathcal{A}(\mathcal{D}_{n};\xi))\geq\epsilon)+\epsilon\mathbb{P}(R_{P}(\mathcal{A}(\mathcal{D}_{n};\xi))<\epsilon)\\ &<B\delta+\epsilon,\end{split}$ (36)
which contradicts with $R_{P,n}(\mathcal{A})\geq\delta B+\epsilon$. This implies that for any $^{\widehat{T}}$ satisfying (8), it also satisfies
$\mathbb{P}_{P}({\widehat{T}}(\mathcal{A},\mathcal{D}_{N})=1)\leq\alpha\textnormal{ for any $\mathcal{A},P$ such that $R_{P,n}(\mathcal{A})\geq\delta B+\epsilon$}.$ (37)
Thus, by Theorem 1, for any $\mathcal{A}$ and any $P$ with $R_{P,n}(\mathcal{A})<\delta B+\epsilon$, the $^{\widehat{T}}$ should satisfy
$\mathbb{P}_{P}({\widehat{T}}(\mathcal{A},\mathcal{D}_{N})=1)\leq\left[\alpha\left(1+\frac{\tilde{\tau}-R_{P,n}(\mathcal{A})}{R_{P}^{\max}-\tilde{\tau}}\right)^{N/n}\right]\wedge 1.$ (38)
Finally, for any $\mathcal{A}$ and any $P$ with $\mathbb{P}(R_{P}(\mathcal{A}(\mathcal{D}_{n};\xi))\geq\epsilon)<\delta$, by (36), we have for the same $\mathcal{A}$ and $P$, it satisfies that $R_{P,n}(\mathcal{A})<B\delta+\epsilon$, then the power upper bound in (38) also holds. This finishes the proof of this proposition.
D.2 Proof of Theorem 2
@@
$\begin{split}&\mathbb{P}({\widehat{T}}(\mathcal{A},\mathcal{D}_{N})=1)\\ &=\mathbb{P}(S<k_{*})+a_{*}\mathbb{P}(S=k_{*})\\ &=(1-a_{*})\mathbb{P}(S<k_{*})+a_{*}\mathbb{P}(S\leq k_{*})\\ &=(1-a_{*})\mathbb{P}(\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},R_{P,n}(\mathcal{A}))<k_{*})+a_{*}\mathbb{P}(\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},R_{P,n}(\mathcal{A}))\leq k_{*})\\ &\leq(1-a_{*})\mathbb{P}(\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},\tau)<k_{*})+a_{*}\mathbb{P}(\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},\tau)\leq k_{*})\\ &=\mathbb{P}(\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},\tau)<k_{*})+a_{*}\mathbb{P}(\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},\tau)=k_{*})\\ &=\alpha\end{split}$
where the last step holds by our choice of $k_{*},a_{*}$, while the inequality holds because $R_{P,n}(\mathcal{A})\geq\tau$ and so $\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},R_{P,n}(\mathcal{A}))$ stochastically dominates $\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},\tau)$.
Next, we compute the power when $\alpha<(1-\tau)^{\lfloor N/(n+1)\rfloor}$. By this bound on $\alpha$ we have
$\mathbb{P}(\textnormal{Binomial}(\textstyle{\lfloor\frac{N}{n+1}\rfloor},\tau)=0)=(1-\tau)^{\lfloor N/(n+1)\rfloor}>\alpha,$
which implies that
@@
Therefore, by the construction of the Binomial test $^{\widehat{T}}$, we have